OpenStudy (anonymous):

FInd the equation of the lines described. Put your final answer in slope-intercept form. Do not graph. slope= 3, y-intercept form (0, -4)

hero (hero):

Hint: Use y = mx + b

OpenStudy (anonymous):

but what is mx+b? i understand thats the slope-intercept form , but how do i use it?

hero (hero):

m = 3 (x,y) = (0,-4) Substitute those values in, then solve for b

OpenStudy (anonymous):

so m is a variable for slope? right? if so then im on track! i understand now thanks,!

hero (hero):

If you understand it, then you should be able to post the slope intercept form of the line.

OpenStudy (anonymous):

this is what i have... y= mx=b -4=3(0) =b -4=0+b *simplifies* -4=b is that right?

OpenStudy (anonymous):

y=mx +b..my bad

hero (hero):

Yes, but you still have to put it in slope intercept form. Basically slope intercept form is where you include m and b values in the formula

hero (hero):

Actually, you can still write the formula and only include m and b: You would get y = 3x -4

OpenStudy (anonymous):

so what would my final answer look like?

OpenStudy (anonymous):

oh lol, okay thanks

hero (hero):

I just showed it to you.